TJ Find the multiplicity of rock 1 f(x) = (x-1)2 inx [EJ Find the order of convergence Pn - ܓܙ 31 10

Please Solve As soon as
Solve quickly I get you thumbs up directly
Thank's
Abdul-Rahim Taysir
Answer #1

solo 야 세 We know ln (1)=0 2 ㅗ Aleksa (3 (1) = 0 20세)와t ) = {() = 0 2 | in + (15) J + 2에 } (-1) ㅅ {: {al + 2 (-1) 부 나 (1) (Y (2) »)- 2 2 lun + belong t 4u-1) - (x-13² x2 Now to 16 (1) + 7.1 = V) - 8 1) = p (1) to Multiplicity of I is 3 T